• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armando Acha

    Log entry time 08:07:17 on May 4,2006

    Entry number 172901

    keyword=Owl Shift Summary

    OWL SHIFT:

    SL : Armando Acha
    TO: Salvatore Frullani

    Evacuated reference cell:


    23:49 Run 4419. Raster 1.5x1.5 Started.
    00:18 Run 4420. Raster 2.0x2.0
    00:35 Run 4421. Raster 2.5x2.5
    00:50 Run 4422. Raster 3.0x3.0

    01:00 Resquested beam off to fill in the cell with H2.

    Target H2:


    01:07 Run 4423. Raster 3.0x3.0.
    01:10 PS2 change to 20 to keep T2 less than 50 Hz.
    01:25 Run 4424. Raster 2.5x2.5.
    01:42 Run 4425. Raster 1.5x1.5.
    02:00 Run 4426. Raster 2.0x2.0.Lack of beam for about 30 min.
    03:00 Stopped run 4426.MCC was doing some beam tunning for Hall B.

    Beam is back.
    03:07 Run 4427. Raster 2.0x2.0. 1M events.


    Target 3He:

    04:05 Target changed: 3He.
    04:08 NMR measurement.
    04:33 BPM Bx and By on target changed to 0.0 and 0.0 (+/- 0.1).
    04:35 Ready to take 15 min runs with different raster sizes on 3He.
    04:36 Parity run 7605 started.
    04:37 Run 4428. Raster 2.0x2.0 (15 min).
    04:56 Run 4429. Raster 3.0x3.0.
    05:15 Run 4430. Raster 2.5x2.5.
    05:42 Run 4431. Raster 1.5x1.5.

    Production (raster 2.0x2.0) :-)

    06:10 Run 4432. Beam current was increased to 4.85 uA, but quickly decreased to 4.2 uA, requested by MCC. Beam not very stable during this run. online replay done
    06:36 Ps factors adjusted to keep T1 and T2 ~ < 50 Hz. New PS1=27500, new PS2 30.

    06:44 Run 4433 started. Online replay done.Detectors are fine.
    07:26 Run 4434 started. Online replay done. Detectors are fine. Abnormal values observed in the Happex feedback monitorat end of run.
    07:57 Happex rebooted.

    Special thanks to Linanga for his help.